if (2006-a)(2004-a)=2005, find (2006-a)^2 + (2004-a)^2.

Answer by ikleyn(52778)

.
From  (2006-a)(2004-a)=2005  you have


    2006*2004 - 2006a - 2004a + a^2 = 2005,

    2006*2004 - 2*2005*a + a^2 = 2005,

    -2*2005*a + a^2 = 2005 - 2006*2004.     (1)


Next,  

     (2006-a)^2 + (2004-a)^2 = 2006^2 -2*2006*a + a^2 + 2004^2 -2*2004*a + a^2 = 

   = 2006^2 + 2004^2 -4*2005a + 2a^2 = 2006^2 + 2004^2 + (2*(-2*2005a  + a^2)) = 


        Now replace the expression  (-2*2005a  + a^2)  by 2005 - 2006*2007, based on (1), and then you get


   = 2006^2 + 2004^2 + 2*(2005 - 2006*2004) = (2006-2004)^2 + 2*2005 = 2^2 + 4010 = 4014.


Answer.  4014.